OpenStudy (anonymous):

Arc length of polar curves. Find arc lenngth of complete circle r = asin(theta) ; a>0.

OpenStudy (anonymous):

Pretty clear on how to find arc length but I think I have the wrong bounds of integration. I got 2pi*a as answer and my book shows pi*a and noted that the bounds of integration where just 0 to pi since that asin(theta) is diameter of circle and that full circle is transversed over interval o to pi.......can someone please help explain this.

myininaya (myininaya):

\[\int\limits_{lower}^{upper} \sqrt{r^2+(\frac{dr}{d \theta})^2} d \theta \\ r=a \sin(\theta)\] so we used the integral right?

myininaya (myininaya):

you should have got pia with those bounds

myininaya (myininaya):

if you used that integral setup

myininaya (myininaya):

you said you used 0=lower and pi=upper?

myininaya (myininaya):

I bet you messed on the trig identity part maybe.. Not totally sure but let's say you did... \[\sqrt{a^2 \sin^2(\theta)+a^2 \cos^2(\theta)}=\sqrt{a^2(\sin^2(\theta)+\cos^2(\theta))} =\sqrt{a^2(1)} \\ =\sqrt{a^2}=a \text{ since } a>0\]

OpenStudy (anonymous):

No I used 0 lower pi upper limits, and the book said to use 0 lower pi upper. I thought since we are in polar to integrate a full circle you would go from 0 to 2pi.

myininaya (myininaya):

oh...no it should be x=0 to pi r=asin(u) u | r 0 a(0)=0 pi/2 | a(1)=a pi | a(0)=0 0 to pi gives that complete rotation

myininaya (myininaya):

because it comes back around to 0

myininaya (myininaya):



myininaya (myininaya):

do you know how graph polor coordinates?

myininaya (myininaya):

polar*

myininaya (myininaya):

if we do theta=0 to theta=2pi it will take us around again

myininaya (myininaya):

u | r 3pi/2, -a